# MonoDGP: Monocular 3D Object Detection with Decoupled-Query and Geometry-Error Priors

### v1

@@ Abstract
Perspective projection has been extensively utilized in monocular 3D object detection methods. It introduces geometric priors from 2D bounding boxes and 3D object dimensions to reduce the uncertainty of depth estimation. However, due to depth errors originating from the object’s visual surface, the height of the bounding box often fails to represent the actual projected central height, which undermines the effectiveness of geometric depth. Direct prediction for the projected height unavoidably results in a loss of 2D priors, while multi-depth prediction with complex branches does not fully leverage geometric depth. This paper presents a Transformer-based monocular 3D object detection method called MonoDGP, which adopts perspective-invariant geometry errors to modify the projection formula. We also try to systematically discuss and explain the mechanisms and efficacy behind geometry errors, which serve as a simple but effective alternative to multi-depth prediction. Additionally, MonoDGP decouples the depth-guided decoder and constructs a 2D decoder only dependent on visual features, providing 2D priors and initializing object queries without the disturbance of 3D detection. To further optimize and fine-tune input tokens of the transformer decoder, we also introduce a Region Segment Head (RSH) that generates enhanced features and segment embeddings. Our monocular method demonstrates state-of-the-art performance on the KITTI benchmark without extra data. Code is available at https://github.com/PuFanqi23/MonoDGP.
@@ 1 Introduction
Figure 1: Comparison with MonoDETR. (b) MonoDGP proposes the RSH for enhanced features and segment embeddings, along with an independent visual decoder for 2D detection and query initialization. (c) Furthermore, we introduce a geometry error prior to convert uneven depth distribution into a more concentrated depth error distribution.
@@
Previous methods [27, 44, 26, 49, 53] typically calculate geometric depth through the projection formula based on bounding boxes and the dimension heights. However, since 2D images are inaccurate in capturing depth errors across an object’s visual surface, the bounding box height is often larger than the projected central height. To address the limitations inherent in single-depth prediction, some approaches[49, 21, 48] integrate multi-depth predictions, including direct depth and geometric depth, and generate final results through weighted averaging. Nevertheless, MonoCD [48] indicates that if the multiple predicted depths cluster on the same side of the ground truth depth, even weighted averaging fails to correct the depth error. DID-M3D [28] decomposes raw depth into visual depth and attribute depth, but ignores geometric depth, using perspective transformation only for data augmentation. Similarly, the geometric depth can be regarded as the visual surface depth, whose value depends on the object’s appearance and position [10]. While the depth error is determined by the inherent attributes of the object and remains invariant with changes in perspective. Besides, the distribution of depth error is more concentrated compared to raw depth. Utilizing error prediction instead of direct prediction is a simple and effective improvement. This approach can not only alleviate the long-tail effect of raw depth distribution, but also avoid the problem where multiple depths fail to complement each other after weighting.
@@
In summary, we propose a transformer-based monocular detector with decoupled queries and geometric errors, called MonoDGP. Our contributions are listed as follows:
  • We adopt perspective-invariant error prediction to replace multi-depth prediction, effectively alleviating the long-tail effect of depth distribution and reducing learning complexities.
@@
  • Experiments illustrate the superiority of MonoDGP. Evaluated on the KITTI 3D object detection benchmark, without any extra data, our method achieves the state-of-the-art (SOTA) performance among monocular detectors.
@@ 2 Related work
Extra training data.  The main challenge of monocular camera-based methods is insufficient of depth cues, which renders 3D object detection an ill-posed problem. Consequently, some approaches [42, 33, 20, 13, 28, 45] incorporate additional data, such as point clouds or depth maps, enabling the model to implicitly learn depth features during training. MonoPGC [45] introduces pixel-level depth map prediction as an auxiliary task, encoding 3D coordinates into depth-perception features to perceive the spatial positions of pixels. CaDDN [33] estimates the classified depth distribution for each pixel, generates high-quality bird’s-eye view feature representations, and achieves accurate 3D object predictions. OccupancyM3D [29] creates occupancy labels from point clouds, and monitors occupancy status in frustum space and 3D space. Given the reliance on expensive sensors and complex collection processes for supplementary information, MonoDGP exclusively utilizes monocular images during the training and inference stages, thereby enhancing the model’s practicality and deployment feasibility.
Transformer-based methods.  Previous works [50, 37, 26, 28, 23, 21] have employed convolutional neural networks (CNNs) to extract local, fine-grained information at various spatial levels. Although CNN-based methods demonstrate superior speed and efficiency, they require multi-layer convolution to progressively aggregate features for capturing long-distance dependencies. Additionally, non-maximum suppression (NMS) technology is employed to filter out redundant detection boxes. Benefiting from the global perception and dynamic adjustment abilities of the attention mechanism [41], an increasing number of studies [13, 49, 52] have shifted their focus toward Transformer-based monocular detectors. For instance, MonoDTR [13] proposes to conduct depth position encoding and integrates global depth information into the Transformer to guide the detection. MonoDETR [49] adheres to the DETR [5] architecture, utilizing object-level depth labels to predict the foreground depth map, constructing a depth-guided decoder, and incorporating object queries for global feature aggregation. MonoDGP also adopts the DETR framework, initializing queries with a decoupled 2D visual decoder and introducing segment embeddings to enhance the contextual relationships among pixels.
@@ 3.1 Overview
The overall framework of MonoDGP is shown in Fig. 2. Firstly, a single-view image is processed by ResNet50 to extract multi-scale feature maps ${f_{i}}\in{\mathbb{R}^{\frac{H}{{{2^{i+3}}}}\times\frac{W}{{{2^{i+3}}}}\times C}},i=0,1,2,3$. The RSH module then predicts target region probabilities and enhances visual features $f_{V}$. A lightweight depth predictor subsequently estimates the depth map and generates depth features $f_{D}$. Both the enhanced visual features $f_{V}$ and depth features $f_{D}$ are fed into two parallel transformer encoders. The transformer decoder consists of a 2D visual decoder and a 3D depth-guided decoder. In the 2D decoder, object queries adaptively achieve semantic scene understanding from visual embeddings, while the 3D decoder extracts geometric and spatial relationships from both visual and depth embeddings. Ultimately, we utilize decoupled queries as inputs for 2D/3D detection heads separately. It is worth noting that after calculating geometric depth based on the projection formula, a geometry error correction mechanism is applied to improve the accuracy of depth estimation.
@@ 3.2 Region Segmentation
Region Segmentation and Enhancement.  We initially utilize a ResNet backbone to downsample and extract multi-scale visual features $f_{0}$ to $f_{3}$. Subsequently, We restore $f_{3}$ to high-resolution feature maps through upsampling operations and skip connections. This integration of shallow and deep features improves multi-scale representation. Specifically, we combine the upsampled $f_{3}$ with $f_{2}$ after a $1\times 1$ convolution and GroupNorm layer. This process is repeated to obtain fused feature maps at various scales. We then employ an SEBlock for attention weighting on each scale’s feature maps. The SEBlock implements a channel attention mechanism through two $1\times 1$ convolution layers: the first convolution reduces the number of feature channels to $1/16$, while the second convolution restores the original number of channels. Finally, we achieve several semantic segmentation maps with the sigmoid function. By multiplying the predicted target probability with original features, RSH effectively enhances foreground features and suppresses background noise.
Figure 4: (a) Schematic illustration of how geometry error occurs. Due to the height of the bounding box $h_{\text{bbox}}$ is larger than the projected central height $h_{c}$, there will be an error between the geometric depth and the ground truth depth. (b) Vehicles at different angles in a bird’s-eye view. The geometric depth $Z_{\text{geo}}$is the distance between the camera plane and the parallel plane on which the closest wheel sits, while the depth error $Z_{\text{err}}$ only depends on dimensions and orientation of the object.
@@
Segment embeddings.  The RSH module generates segment embeddings from target region probabilities that exceed a threshold of 0.5. We subsequently add these segment embeddings to the token and position embeddings. This approach is inspired by the segment embeddings mechanism used in the BERT model. Segment embeddings help to better distinguish the semantic relationships among image tokens, similar to how different “sentences” are represented in BERT.
@@ 3.3 Decoupled Query
After obtaining visual embeddings $f_{V}^{e}$ and depth embeddings $f_{D}^{e}$ from two parallel encoders, MonoDGP decouples the depth-guided decoder into a 2D decoder and a 3D decoder. By initializing queries and reference points with the 2D visual decoder, the object queries can acquire spatial position, shape and category information, reducing depth uncertainties from geometric priors. These 2D queries are then fed into the 3D decoder to interact with visual and depth features, allowing each decoder to focus on its own specific task and ensuring stable convergence during training.
@@
3D detection decoder.  The initialized 2D object queries ${q_{2d}}\in{R^{N\times C}}$ are input to the 3D depth-guided decoder. Each decoder block contains a depth cross-attention layer, an inter-query self-attention layer, a visual cross-attention layer and an FFN. Specifically, object queries first extract geometric and perspective cues from depth embeddings $f_{D}^{e}$. Subsequently, the depth-aware queries are sent to the inter-query self-attention layer to promote information exchange. After the visual cross-attention layer and FFN, we generate final 3D queries ${q_{3d}}$ that combine semantic and geometric information, obtaining a more accurate scene-level spatial understanding.
@@ 3.4 Geometry Error
Direct depth estimation is often ill-posed in monocular methods. We apply the perspective projection formula $Z=f*H_{3D}/h_{2D}$ to introduce geometric priors and mitigate depth uncertainty. When utilizing the projection formula, it is essential to make sure that the object’s dimension height (such as $AB$ and $CD$ in Fig. 4(a)) is parallel to the projection plane, namely the X-Y plane in the camera coordinate system, and remains perpendicular to the X-Z plane. To guarantee the validity of the formula, we establish the following assumptions:
@@
The perspective projection formula contains two variables: the height of the 2D bounding box and the height of the 3D size. To correct the formula, the depth error is added to the geometric depth for the final prediction, while the errors from both heights can also compensate for the geometric depth. Their error predictions are formulated as:
@@
Test, $AP_{BEV}$ Test, $AP_{3D}$ Val, $AP_{BEV}$ Val, $AP_{3D}$
Method Extra data Reference Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard
@@
Improvement - - -0.14 +1.05 +0.65 +0.80 +1.55 +1.18 +1.54 +1.25 +1.33 +1.92 +1.73 +1.87
3.5 Why Geometry Error Works?
  • Different geometry errors: why we select depth error?
2D and 3D attributes.  The generation of bounding boxes relies on 2D visual features like color, edges, textures, and shapes. In contrast, predicting 3D attributes requires extra factors including depth positioning and geometric perspective, which are difficult to obtain from single-view images. Benefiting from the network’s powerful capability of marginal feature extraction, $h_{bbox}$ can provide more accurate 2D priors. However, we can hardly predict $h_{c}$ from direct marginal features. The geometric depth calculated by $h_{bbox}$ should be regarded as a reliable surface depth, rather than an inaccurate central depth.
Different 3D attributes.  In the absence of direct depth information, the 3D data distribution will determine the learning complexity of the network. The Fig. 5 and 6 show the data distributions and boxplot of depth-related attributes. Based on these figures, we can further evaluate and compare the learning difficulty across different data distributions. The central depth usually exhibits a long-tailed distribution characterized by a wide and uneven range. Conversely, the 3D size distribution tends to be more concentrated. For instance, most cars are around 1.5 meters tall, which facilitates the model to learn and discern the underlying patterns within the data distribution. The boxplot indicates that the depth distribution has a broader interquartile range (IQR) and longer whiskers, suggesting greater complexity and noise. Meanwhile, the IQR for dimension height, depth error, and height error is narrower, which illustrates that these distributions are relatively compact with plain patterns.
@@ 3.6 Loss Function
The training loss consists of four distinct components: region loss $L_{region}$, depth map loss $L_{dmap}$, 2D loss $L_{2D}$, and 3D loss $L_{3D}$. Specifically, $L_{2D}$ contains the losses of object category, projected center point, and bounding box, while $L_{3D}$ includes the losses of orientation, 3D size, and central depth. We denote the weights of these losses from ${\lambda_{1}}$ to ${\lambda_{6}}$. The overall loss of MonoDGP is formulated as:
${L_{overall}}={L_{2D}}+{L_{3D}}+{\lambda_{7}}{L_{dmap}}+{\lambda_{8}}\sum\limits_{i=0}^{4}{L_{region}^{i}}$ (10)
@@ 4.1 Settings
Evaluation Metrics.  We present the detection outcomes categorized by three difficulty levels: easy, moderate, and hard. Consistent with prior works, our evaluation relies on average precision (AP) metrics for both 3D bounding boxes and bird’s-eye view, referred to as $AP_{3D}$ and $AP_{BEV}$, which are calculated at 40 recall positions following the established protocol [39]. The benchmark ranks all methods based on the moderate $AP_{3D}$ metrics of "Car" category. An IoU threshold of 0.7 is also applied during evaluation.
Implementation Details.  We take ResNet50 [12] as our backbone. In regard to Transformer, we adopt 8 heads in each attention module and set the number of queries 50 with 4 sampling points in the multi-scale deformable attention. The weights of losses are set as {5, 2, 10, 1, 1, 1, 1, 1} for ${\lambda_{1}}$ to ${\lambda_{8}}$. On a single RTX 3090 GPU, using Mixup3D, we train MonoDGP for 250 epochs with a batch size of 8 and a learning rate of 2e-4. We utilize AdamW [25] optimizer with weight decay and decrease learning rate by a factor of 0.5 at epochs 85, 125, 165, and 225. To accelerate the model’s convergence, we adopt group-wise one-to-many assignment [6] with 11 groups. During inference, we filter out queries with category confidence lower than 0.2.

### v2

@@ Abstract
Perspective projection has been extensively utilized in monocular 3D object detection methods. It introduces geometric priors from 2D bounding boxes and 3D object dimensions to reduce the uncertainty of depth estimation. However, due to errors originating from the object’s visual surface, the bounding box height often fails to represent the actual central height, which undermines the effectiveness of geometric depth. Direct prediction for the projected height unavoidably results in a loss of 2D priors, while multi-depth prediction with complex branches does not fully leverage geometric depth. This paper presents a Transformer-based monocular 3D object detection method called MonoDGP, which adopts perspective-invariant geometry errors to modify the projection formula. We also try to systematically discuss and explain the mechanisms and efficacy behind geometry errors, which serve as a simple but effective alternative to multi-depth prediction. Additionally, MonoDGP decouples the depth-guided decoder and constructs a 2D decoder only dependent on visual features, providing 2D priors and initializing object queries without the disturbance of 3D detection. To further optimize and fine-tune input tokens of the transformer decoder, we also introduce a Region Segmentation Head (RSH) that generates enhanced features and segment embeddings. Our monocular method demonstrates state-of-the-art performance on the KITTI benchmark without extra data. Code is available at https://github.com/PuFanqi23/MonoDGP.
@@ 1 Introduction
Figure 1: Comparison with MonoDETR. MonoDGP employs an RSH module for enhanced features and segment embeddings, along with an independent visual decoder for 2D query initialization. We further propose a geometry error prior that converts uneven depth distribution into more concentrated error distribution.
@@
Previous methods like Deep3DBox [36] and Shift R-CNN [37] enforce geometric constraints through tight alignment between 3D box projections and 2D detections. Unlike strict constraints mentioned above, geometric depth is uniquely determined by the ratio of an object’s 3D height to its 2D projected height. This physically interpretable explicit mapping eliminates cascaded uncertainty, ensuring superior computational efficiency and noise robustness. However, since 2D images are inaccurate in capturing depth errors of the object visual surface, the 2D bounding box height often exceeds the projected height, which makes geometric depth deviate from the actual depth. To address single-depth prediction limitations, some approaches[62, 26, 60] integrate multi-depth predictions through weighted averaging, including direct and geometric depths. Nevertheless, MonoCD [60] indicates that if these predicted depths cluster on the same side of the ground truth depth, even weighted averaging fails to correct the depth error. DID-M3D [38] decomposes depth into visual and attribute parts, using perspective transformation for data augmentation. Similarly, the geometric depth can also be regarded as the visual depth, whose value depends on the object’s surface appearance and position [13]. While the depth error is determined by the object’s inherent attributes and remains invariant with changes in perspective. With more concentrated distribution, error prediction can not only alleviate the long-tail effect of raw depth, but also avoid the problem where multiple depths fail to complement each other after weighting.
@@
In summary, we propose a transformer-based method called MonoDGP. Our contributions are listed as follows:
  • We adopt perspective-invariant error prediction to replace multi-depth prediction, alleviating the long-tail effect of depth distribution and reducing learning complexities.
@@
  • Evaluated on the KITTI 3D object detection benchmark, without any extra data, MonoDGP achieves the state-of-the-art (SOTA) performance among monocular detectors.
@@ 2 Related work
Extra training data.  The main challenge of monocular methods is insufficient of depth cues, which renders 3D object detection an ill-posed problem. Consequently, some approaches [54, 43, 25, 17, 38, 57] incorporate additional data, such as point clouds or depth maps, enabling the model to implicitly learn depth features during training. MonoPGC [57] introduces pixel-level depth map prediction as an auxiliary task, encoding 3D coordinates into depth-perception features to perceive the spatial positions of pixels. CaDDN [43] estimates the classified depth distribution for each pixel, generates high-quality bird’s-eye view feature representations, and achieves accurate 3D object predictions. OccupancyM3D [39] creates occupancy labels from point clouds, and monitors occupancy status in frustum space and 3D space. Given the reliance on expensive sensors and complex collection processes for supplementary information, MonoDGP exclusively utilizes monocular images during the training and inference stages, thereby enhancing the model’s practicality and deployment feasibility.
Transformer-based methods.  Many previous works [63, 48, 33, 38, 30, 26] have employed convolutional neural networks (CNNs) to extract local, fine-grained information at various spatial levels. Although CNN-based methods demonstrate superior speed and efficiency, they require multi-layer convolution to progressively aggregate features for capturing long-distance dependencies. Additionally, non-maximum suppression (NMS) technology is employed to filter out redundant detection boxes. Benefiting from the global perception and dynamic adjustment abilities of the attention mechanism [53], an increasing number of studies [17, 62, 65] have shifted their focus toward Transformer-based monocular detectors. For instance, MonoDTR [17] proposes to conduct depth position encoding and integrates global depth information into the Transformer to guide the detection. MonoDETR [62] adheres to the DETR [8] architecture, utilizing object-level depth labels to predict the foreground depth map, constructing a depth-guided decoder, and incorporating object queries for global feature aggregation. MonoDGP also adopts the DETR framework, initializing queries with a decoupled 2D visual decoder and introducing segment embeddings to enhance the contextual relationships among pixels.
@@ 3.1 Overview
The overall framework of MonoDGP is shown in Fig. 2. Firstly, a single-view image is processed by ResNet50 to extract multi-scale feature maps ${f_{i}}\in{\mathbb{R}^{\frac{H}{{{2^{i+3}}}}\times\frac{W}{{{2^{i+3}}}}\times C}},i=0,1,2,3$. The RSH module then predicts target region probabilities and enhances visual features $f_{V}$. A lightweight depth predictor subsequently estimates the depth map and generates depth features $f_{D}$. Both enhanced visual features $f_{V}$ and depth features $f_{D}$ are fed into two parallel transformer encoders. The transformer decoder consists of a 2D visual decoder and a 3D depth-guided decoder. In the 2D decoder, object queries adaptively achieve semantic scene understanding from visual embeddings, while the 3D decoder extracts geometric and spatial relationships from both visual and depth embeddings. Ultimately, we utilize decoupled queries as inputs for 2D/3D detection heads separately. It is worth noting that after calculating geometric depth based on the projection formula, a geometry error correction mechanism is applied to improve the accuracy of depth estimation.
@@ 3.2 Region Segmentation
Region Segmentation and Enhancement.  We initially utilize a ResNet backbone to downsample and extract multi-scale visual features $f_{0}$ to $f_{3}$. Subsequently, We restore $f_{3}$ to high-resolution feature maps through upsampling operations and skip connections. This integration of shallow and deep features improves multi-scale representation. Specifically, we combine the upsampled $f_{3}$ with $f_{2}$ after a $1\times 1$ convolution and GroupNorm layer. This process is repeated to obtain fused feature maps at various scales. We then employ an SEBlock [16] for attention weighting on each scale’s feature maps. The SEBlock implements a channel attention mechanism through two $1\times 1$ convolution layers: the first convolution reduces the number of feature channels to $1/16$, while the second convolution restores the original number of channels. Finally, we achieve several semantic segmentation maps with the sigmoid function. By multiplying the predicted target probability with original features, RSH effectively enhances foreground features and suppresses background noise.
Figure 4: (a) Schematic illustration of how geometry error occurs. Due to the height of the bounding box $h_{\text{bbox}}$ is larger than the projected central height $h_{c}$, there will be an error between geometric depth and ground truth depth. (b) Vehicles at different angles in a bird’s-eye view. The geometric depth $Z_{\text{geo}}$ is the distance between the camera plane and the parallel plane on which the closest wheel sits, while depth error $Z_{\text{err}}$ only depends on the object’s dimensions and orientation. The camera height also has a tiny effect on the $Z_{\text{err}}$ , rigorous proof is given in the appendix.
@@
Segment embeddings.  The RSH module generates segment embeddings from target region probabilities that exceed a threshold of 0.5. We subsequently add these segment embeddings to the token and position embeddings. This approach is inspired by the segment embeddings mechanism utilized in the BERT model [19]. Segment embeddings help to better distinguish the semantic relationships between foreground and background image tokens, similar to how different “sentences” are represented in BERT.
@@ 3.3 Decoupled Query
Geometric priors optimize depth estimation, but also introduce depth uncertainty to 2D detection. To mitigate this side effect, we decouple object queries for independent 2D detection. After obtaining visual embeddings $f_{V}^{e}$ and depth embeddings $f_{D}^{e}$ from two parallel encoders, MonoDGP decomposes the depth-guided decoder into a 2D decoder and a 3D decoder. The 2D visual decoder initialize 2D queries and reference points to acquire spatial position and category information. These 2D queries are then fed into the 3D decoder to interact with visual and depth features. This allows each decoder to focus on its own specific task and ensures stable convergence during training.
@@
3D detection decoder.  The initialized 2D object queries ${q_{2d}}\in{R^{N\times C}}$ are input to the 3D depth-guided decoder. Each decoder block contains a depth cross-attention layer, an inter-query self-attention layer, a visual cross-attention layer and an FFN. Specifically, object queries first extract perspective cues from depth embeddings $f_{D}^{e}$. Subsequently, the depth-aware queries are sent to the inter-query self-attention layer to promote information exchange. After the visual cross-attention layer and FFN, we generate 3D queries ${q_{3d}}$ that combine semantic and geometric features, obtaining an accurate scene-level spatial understanding.
@@ 3.4 Geometry Error
Direct depth estimation is often ill-posed in monocular methods. We apply the perspective projection formula $Z=f\times H_{3D}/h_{2D}$ to introduce geometric priors and mitigate depth uncertainty. When utilizing the projection formula, it is essential to make sure that the object’s dimension height (such as $AB$ and $CD$ in Fig. 4(a)) is parallel to the projection plane, namely the X-Y plane in the camera coordinate system, and remains perpendicular to the X-Z plane. To guarantee the validity of the formula, we establish the following assumptions:
@@
The perspective projection formula contains two variables: the height of the 2D bounding box and the height of the 3D size. To correct the formula, the depth error is added to the geometric depth for the final prediction, while errors from both the 2D and 3D heights can also compensate for the geometric depth. Their predictions are formulated as:
@@
3.5 Why Geometry Error Works?
  • Different geometry errors: why we select depth error?
2D and 3D attributes.  2D predictions rely on monocular visual features like color, edges, textures, whereas 3D predictions require extra factors including depth positioning and geometric perspective, which are difficult to obtain from single-view images. Benefiting from the network’s powerful capability of marginal feature extraction, $h_{bbox}$ can provide more accurate 2D priors. However, we can hardly predict $h_{c}$ from direct marginal features. The geometric depth calculated by $h_{bbox}$ should be regarded as a reliable surface depth, rather than an inaccurate central depth.
Test, $AP_{BEV|R40}$ Test, $AP_{3D|R40}$ Val, $AP_{BEV|R40}$ Val, $AP_{3D|R40}$
Methods Extra data Reference Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard
@@ 3.5 Why Geometry Error Works?
Improvement - - -0.14 +1.05 +0.65 +0.80 +1.55 +0.87 +1.54 +1.25 +1.33 +1.92 +1.73 +1.87
Different 3D attributes.  Without direct depth supervision, data distribution dictates network learning complexity. The Fig. 5 and 6 show the data distributions and boxplot of depth-related 3D attributes. Based on these figures, we can further evaluate and compare the learning difficulty across different data distributions. The central depth exhibits a long-tailed distribution with wide variance, while the 3D size distribution tends to be more concentrated. For instance, typical car height clusters near 1.5 meters, simplifying pattern recognition. The boxplot indicates that the depth distribution has a broader interquartile range (IQR) and longer whiskers, suggesting greater complexity and noise. Meanwhile, the IQR for dimension height, depth error, and height error is narrower, which illustrates that these distributions are relatively compact with plain patterns.
@@ 3.6 Loss Function
The training loss consists of four distinct components: region loss $L_{region}$, depth map loss $L_{dmap}$, 2D loss $L_{2D}$, and 3D loss $L_{3D}$. Specifically, $L_{2D}$ contains the losses of object category, bounding box, GIoU, and projected center point, while $L_{3D}$ includes the losses of 3D size, orientation, and central depth. We denote the weights of 2D and 3D losses from ${\lambda_{1}}$ to ${\lambda_{7}}$. The overall loss is formulated as:
${L_{overall}}={L_{2D}}+{L_{3D}}+{\lambda_{8}}{L_{dmap}}+{\lambda_{9}}\sum\limits_{i=0}^{4}{L_{region}^{i}}$ (10)
@@ 4.1 Settings
Evaluation Metrics.  We present the detection outcomes categorized by three difficulty levels: easy, moderate, and hard. Consistent with prior works, our evaluation relies on average precision (AP) metrics for both 3D bounding boxes and bird’s-eye view, referred to as $AP_{3D}$ and $AP_{BEV}$, which are calculated at 40 recall positions following the established protocol [50]. The benchmark ranks all methods based on the moderate $AP_{3D}$ metrics of “Car” category.
Implementation Details.  We take ResNet50 [15] as our backbone. In regard to Transformer, we adopt 8 heads in each attention module and set the number of queries 50 with 4 sampling points in the multi-scale deformable attention. The weights of losses are set as {2, 5, 2, 10, 1, 1, 1, 1, 1} for ${\lambda_{1}}$ to ${\lambda_{9}}$. On a single RTX 3090 GPU, using Mixup3D [27], we train MonoDGP for 250 epochs with a batch size of 8 and a learning rate of 2e-4. We utilize AdamW [32] optimizer with weight decay and decrease learning rate by a factor of 0.5 at epochs 85, 125, 165, and 225. To accelerate the model’s convergence, we adopt group-wise one-to-many assignment [9] with 11 groups. During inference, we filter out queries with category confidence lower than 0.2.
Methods Val, Mod. $AP_{3D}$ Params (M)$\downarrow$ FLOPs (G)$\downarrow$ Runtime (ms)$\downarrow$
Baseline 20.22 35.93 59.72 35
MonoDGP (Ours) 22.34 38.90 68.99 42
w/o Decoupled Query 21.48 37.54 65.75 39
w/o RSH 21.93 36.23 62.25 37
Table 2: Ablation study of computational cost. We test the Runtime (ms) on a single RTX 3090 GPU with a batch size of 1.